an ideal refrigerator runs etween – 23 C and 27 C find heat rejected to atmosphere for every joule of work
 

Grade:12

1 Answers

Khimraj
6 years ago
Let heat rejected is Q1 = x and W = 1J.
Then Q2 = Q1 – W = x-1
T1 = 27 C= 300 K.
T2 = -23 C = 250 K
For ideal process Q2/Q1 = T2/T1
x -1/x = 300/250
So x = 5J